A sample of gas has a mass of 0.530 g . Its volume is 115 mL at a temperature of 75 ∘C and a pressure of 762 mmHg .
1
Expert's answer
2017-04-03T10:55:07-0400
762 mmHg x (1 atm / 760 mmHg) = 1.002 atm
PV = nRT
n = PV/RT
n = (1.002 atm x 0.115 L) / (0.08206 L atm / K mol x 348 K) = 0.004035 mol
Molar mass = g/mol = 0.530 g / 0.004035 mol = 131 g/mol


Answer: M = 131 g/mol